# Photoemission tomography of excitons in 2D systems: momentum-space signatures of correlated electron-hole wave functions

### v1

@@ III.3 Pump polarization effects in 2D hBN
Figure 6: Influence of pump laser’s polarization on exciton photoemission. The four columns shows ARPES photoemission maps for different polarizations of the pump laser (black arrows). (a) shows photoemission from the lowest-energy exciton while (b) shows the second exciton. In the case of the first exciton, only small differences in the intensity distribution along the horseshoes is visible, while for the second exciton, there is a strong dependence on the pump polarization. Different polarizations lead to different rotations of the horseshoes. In the case of both excitons, the photoemission maps are mirror symmetric across the (high-symmetry) polarization directions.
As outlined in Paragraph II.2, the summation of energy-degenerate BSE eigenvectors to form a superimposed exciton involves coefficients that are generally non-unique and depend on the polarization of the pump pulse. The influence of pump polarization is illustrated in detail in Figure 6. Panels (a) and (b) in Figure 6 depict the photoemission from the $S_{1}$ exciton (5.3 eV) at a kinetic energy of 23.9 eV and the $S_{2}$ exciton (6.2 eV) at 24.8 eV, respectively. The black arrows indicate the polarization direction of the pump-pulse. Photoemission from the first exciton is only weakly dependent on the pump-pulse’s polarization and retains the horseshoe intensity pattern. Varying the polarization reveals minor variations in the intensity distributions along the horseshoes. For the second exciton, however, there is a strong dependence of the photoemission intensity distribution on the pump pulse, where different polarizations lead to different redistribution of spectral weight within the horseshoes, allowing for precise control of this redistribution based on the polarization direction. Although the probe pulse polarization may introduce an overall attenuation in an trARPES experiment, pump-induced effects on exciton photoemission are expected to be observable, provided that sufficient energy and momentum resolution are achieved.
@@ IV Conclusions
We have developed a framework for exciton photoemission orbital tomography (exPOT) in periodic systems, thereby enabling the computation and interpretation of photoemission angular distributions from excitons in crystalline two-dimensional materials. The exPOT formalism is derived from first principles, using many-body perturbation theory and the Dyson orbital concept, and connects the photoemission intensity to results from GWGW+BSE calculations. As a manifestation of the electron-hole correlation, for each hole configuration, exciton photoemission is expressed as a coherent sum over Fourier-transformed conduction-band wave functions, weighted by the Bethe–Salpeter eigenvectors.
@@
The exPOT formalism for periodic systems bridges the gap between ab-initio many-body theory and experimentally measurable momentum-resolved spectra of excitons. It provides a powerful tool for gaining microscopic insight into the internal structure of excitons and offers a predictive framework for designing and interpreting next-generation trARPES experiments on low-dimensional quantum materials. Our formalism also includes excitons with finite center-of-mass momentum, making it particularly suited for elucidating the nature of momentum-dark excitons by establishing a direct connection between theory and experiment.
@@ Appendix A Expansion of Dyson orbitals in Bloch states
In this section, we derive the expansion in Equation II.1 of the Dyson orbitals, as described in Equation II.1, into Bloch functions.
Since this expansion holds in general, we start with the Dyson orbital formed by the overlap of an arbitrary N1N-1- and $N$-electron wave function:
@@
$\Psi^{N}\left(\boldsymbol{r},\boldsymbol{r}_{2},...,\boldsymbol{r}_{N}\right)$ (21)
Next, since the Dyson-orbitals are one-electron wave functions, they can be expanded in any set of one-electron basis functions. For a set of (BvK-periodic, just like $D$) orthonormal basis functions $\{\psi_{j}(\cdot)\}$ we can thus write
@@
$c_{j}=\int_{V}d^{3}r\psi_{j}(\boldsymbol{r})^{*}D(\boldsymbol{r}).$ (23)
Inserting the Dyson orbital and renaming $\boldsymbol{r}$ to $\boldsymbol{r}_{1}$, we get
@@
where we used the Pauli exclusion principle in the last step.
Using the same argument, we find for any $i$ in $\{1,...,N\}$:
@@
$\Psi^{N-1}\left(\boldsymbol{r}_{1},...,\boldsymbol{r}_{i-1},\boldsymbol{r}_{i+1},...,\boldsymbol{r}_{N}\right)^{*}\Psi^{N}\left(\boldsymbol{r}_{1},\boldsymbol{r}_{2},...,\boldsymbol{r}_{N}\right)$ (26)
@@
Now, note, that the expression in parenthesis is precisely the antisymmetrized product $|{\mathcal{A}}\psi_{j}\Psi^{N-1}\rangle$ (see, e.g., [37]). Using the creation operator $\hat{a}_{j}^{\dagger}$ corresponding to $|\psi_{j}\rangle$ and a bra-ket notation, we can thus write
@@
Inserting this into Equation 22, we get our final result for the expansion of the Dyson orbital $D$ in an arbitrary, orthonormal basis set $\{\psi_{j}(\cdot)\}$:
$D(\boldsymbol{r})=$ $\sum_{j}\langle\Psi^{N-1}|\hat{a}_{j}|\Psi^{N}\rangle\psi_{j}(\boldsymbol{r})$ (33)
@@
To get to Equation II.1, we then simply use the Bloch functions as basis functions and split the sum over bands into a sum over valence bands and a sum over conduction bands.
Appendix B Computational Details
@@
For the example of monolayer hBN, we calculate Kohn-Sham DFT wave functions and energies in Quantum Espresso [42, 43, 44] using the PBE exchange-correlation functional, ONCV pseudopotentials and a cutoff energy of 50 Ry and 200 Ry for the wave functions and the representation of the density, respectively. The lattice constant was set to 2.507 $\mathrm{\AA }$ (bulk experimental value: 2.506 $\mathrm{\AA }$ [54]) and we use a vacuum layer of 20 $\mathrm{\AA }$, as well as a truncation of the Coulomb interaction in $z$-direction [55], in order to avoid spurious interaction between repeated slabs from the periodic boundary condition. Based on these ground-state properties, we perform a $G_{0}W_{0}$ calculation using BerkeleyGW [1, 2, 46]. In particular, we use the NNS subsampling technique, which, for systems with reduced dimensionality, accelerates convergence with respect to the Brillouin zone sampling [56]. For the calculation of the static inverse dielectric function we use a 12x12x1 sampling of the Brillouin zone, a cutoff energy of 30 Ry for reciprocal lattice vectors and sum over 200 bands. Based on this, we calculate the $G_{0}W_{0}$ energies for 4 valence and 44 conduction bands, using the same k-space mesh and reciprocal lattice cutoff as for the inverse dielectric function.
@@
Finally, we construct the BSE Hamiltonian using 4 valence and 9 conduction bands on a 12x12x1 k-space mesh, interpolate it to a 96x96x1 k-space mesh (4 valence, 4 conduction bands) and diagonalize the BSE Hamiltonian.

### v2

@@ III.3 Pump polarization effects in 2D hBN
Figure 6: Influence of pump laser’s polarization on exciton photoemission. The four columns shows ARPES photoemission maps for different polarizations of the pump laser (black arrows). (a) shows photoemission from the lowest-energy exciton while (b) shows the second exciton. In the case of the first exciton, only small differences in the intensity distribution along the horseshoes is visible, while for the second exciton, there is a strong dependence on the pump polarization, with different polarizations leading to different rotations of the horseshoes. In the case of both excitons, the photoemission maps are mirror symmetric across the (high-symmetry) polarization directions.
As outlined in Sec. II.2, the summation of energy-degenerate BSE eigenvectors, in order to form a superimposed exciton, involves coefficients that depend on the polarization of the pump pulse. The influence of pump polarization is illustrated in detail in Fig. 6. Panels (a) and (b) in Fig. 6 depict the photoemission from the $S_{1}$ exciton (5.3 eV) at a kinetic energy of 23.9 eV and the $S_{2}$ exciton (6.2 eV) at 24.8 eV, respectively. The black arrows indicate the polarization direction of the pump pulse. Photoemission from the first exciton is only weakly dependent on the pump pulse’s polarization and retains the horseshoe intensity pattern. Varying the polarization reveals minor variations in the intensity distributions along the horseshoes. For the second exciton, however, there is a strong dependence of the photoemission intensity distribution on the pump pulse, where different polarization directions lead to different redistributions of spectral weight within the horseshoes. Although the probe pulse polarization may introduce an overall attenuation in an trARPES experiment, pump-induced effects on exciton photoemission are therefore expected to be observable, provided that sufficient energy and momentum resolution are achieved.
@@ IV Conclusions
We have developed a framework for exciton photoemission orbital tomography (exPOT) in periodic systems, thereby enabling the computation and interpretation of photoemission angular distributions from excitons in crystalline two-dimensional materials. The exPOT formalism is derived from first principles, using many-body perturbation theory and the Dyson orbital concept, and connects the photoemission intensity to results from GWGW+Bethe-Salpeter calculations. As a manifestation of the electron-hole correlation, for each hole configuration, exciton photoemission is expressed as a coherent sum over Fourier-transformed conduction-band wave functions, weighted by the Bethe–Salpeter eigenvector components.
@@
The exPOT formalism for periodic systems provides a powerful tool for gaining microscopic insight into the internal structure of excitons and offers a predictive framework for designing and interpreting next-generation trARPES experiments on quantum materials. Our formalism also includes excitons with finite center-of-mass momentum, making it particularly suited for elucidating the nature of momentum-dark excitons by bridging the gap between ab-initio many-body theory and experimentally measurable momentum-resolved spectra.
@@ Appendix A Expansion of Dyson orbitals in Bloch states
In this section, we derive the expansion Eq. II.1 of the Dyson orbitals, as given in Eq. II.1, in terms of Bloch functions.
Since this expansion holds in general, we start with the Dyson orbital formed by the overlap of an arbitrary (N1)(N-1)- and $N$-electron wave function:
@@
$\Psi^{N}\left(\boldsymbol{r},\boldsymbol{r}_{2},...,\boldsymbol{r}_{N}\right).$ (20)
Next, since the Dyson-orbitals are one-electron wave functions, they can be expanded in any set of one-electron basis functions. For a set of (Born-von Kármán-periodic, BvK-periodic, just like $D$) orthonormal basis functions $\{\psi_{j}(\cdot)\}$ we can thus write
@@
$c_{j}=\int_{V}d^{3}r\psi_{j}(\boldsymbol{r})^{*}D(\boldsymbol{r}),$ (22)
where $V$ is the crystal volume, see text below Eq. II.1. Inserting the Dyson orbital and renaming $\boldsymbol{r}$ to $\boldsymbol{r}_{1}$, we get
@@
where we used the Pauli exclusion principle in the last step. Using the same argument, we find for any $i$ in $\{1,...,N\}$:
@@
$\Psi^{N-1}\left(\boldsymbol{r}_{1},...,\boldsymbol{r}_{i-1},\boldsymbol{r}_{i+1},...,\boldsymbol{r}_{N}\right)^{*}\Psi^{N}\left(\boldsymbol{r}_{1},\boldsymbol{r}_{2},...,\boldsymbol{r}_{N}\right).$ (25)
@@
Now we can identify the expression in parentheses with the wave function of the anti-symmetrized product state $|{\mathcal{A}}\psi_{j}\Psi^{N-1}\rangle$ (see, e.g., Ref. [37]). Using the creation operator $\hat{a}_{j}^{\dagger}$ corresponding to $|\psi_{j}\rangle$ and a bra-ket notation, we can thus write
@@
Inserting this into Eq. 21, we get our final result for the expansion of the Dyson orbital $D$ in an arbitrary, orthonormal basis set $\{\psi_{j}(\cdot)\}$:
$D(\boldsymbol{r})=$ $\sum_{j}\langle\Psi^{N-1}|\hat{a}_{j}|\Psi^{N}\rangle\psi_{j}(\boldsymbol{r}).$ (32)
@@
To get to Eq. II.1, we then simply use the Bloch functions as basis functions and split the sum over bands into a sum over valence bands and a sum over conduction bands.
Appendix B Implementation Details
This section provides Eq. II.1 in an alternative form that allows for efficient implementation when using a plane-wave basis set. First note, that in Eq. II.1 in the Brillouin-zone (BZ) summation over the Bloch momentum $\boldsymbol{q}$, only a single $\boldsymbol{q}$ vector contributes. This can be seen, without loss of generality, by expanding $\xi_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{r})$ in a plane-wave basis (using Born-von Kármán, BvK, periodic boundary conditions):
$\xi_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{r})$ $=\frac{1}{V}\sum_{\boldsymbol{G}}c_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{G})e^{\mathrm{i}(\boldsymbol{q}+\boldsymbol{Q}+\boldsymbol{G})\boldsymbol{r}},$
$c_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{G})$ $=\int_{V}\mathrm{d}^{3}r\xi_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{r})e^{-\mathrm{i}(\boldsymbol{q}+\boldsymbol{Q}+\boldsymbol{G})\boldsymbol{r}}.$ (33)
The Fourier-transform
$\mathcal{F}[\xi_{c,\boldsymbol{q}+\boldsymbol{Q}}](\boldsymbol{k})=\int_{V}\mathrm{d}^{3}r\xi_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{r})e^{-\mathrm{i}\boldsymbol{k}\boldsymbol{r}}$ (34)
then vanishes unless $\boldsymbol{q}+\boldsymbol{G}+\boldsymbol{Q}=\boldsymbol{k}$ for at least one of the reciprocal lattice vectors $\boldsymbol{G}$. Now, since $\boldsymbol{q}$ is restricted to the BZ and $\boldsymbol{G}$ is a reciprocal lattice vector, there is at most a single combination of $\boldsymbol{q}$ and $\boldsymbol{G}$ that satisfies this condition. Thus, the BZ summation is reduced to a single contributing $\boldsymbol{q}$ vector and
$\mathcal{F}[\xi_{c,\boldsymbol{q}+\boldsymbol{Q}}](\boldsymbol{k})=\sum_{\boldsymbol{G}}c_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{G})\delta(\boldsymbol{q}+\boldsymbol{G}+\boldsymbol{Q}-\boldsymbol{k}).$ (35)
The photoemission intensity may then be written as
$I_{m,\boldsymbol{Q}}(\boldsymbol{k})\propto|\boldsymbol{A}\boldsymbol{k}|^{2}\sum_{v}\sum_{\boldsymbol{q}}^{\mathrm{BZ}}\sum_{\boldsymbol{G}}\delta(\boldsymbol{q}+\boldsymbol{G}+\boldsymbol{Q}-\boldsymbol{k})$
$\cdot\left|\sum_{c}X_{v,c,\boldsymbol{q}}^{(m,\boldsymbol{Q})}c_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{G})\right|^{2}\delta\left(\omega+\Omega^{(m,\boldsymbol{Q})}-\varepsilon_{v,\boldsymbol{q}-E_{\mathrm{kin}}}\right).$ (36)
This can be used for an efficient implementation of the presented formalism: The grid of $\boldsymbol{G}$ vectors, determined by the used unit cell and the grid of $\boldsymbol{q}$ vectors, determined by the used BZ sampling, create a grid of $\boldsymbol{k}$ vectors, $\boldsymbol{k}=\boldsymbol{q}+\boldsymbol{G}+\boldsymbol{Q}$, for which we get:
$I_{m,\boldsymbol{Q}}(\boldsymbol{k})\propto$ $|\boldsymbol{A}\boldsymbol{k}|^{2}\sum_{v}\left|\sum_{c}X_{v,c,\boldsymbol{q}}^{(m,\boldsymbol{Q})}c_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{G})\right|^{2}$
$\cdot\delta\left(\omega+\Omega^{(m,\boldsymbol{Q})}-\varepsilon_{v,\boldsymbol{q}}-E_{\mathrm{kin}}\right),$
$\mathrm{with}~\boldsymbol{q}+\boldsymbol{G}+\boldsymbol{Q}=\boldsymbol{k}.$ (37)
For any photoelectron momentum $\boldsymbol{k}$ on this grid, we can thus directly use the plane-wave coefficients $c_{c,\boldsymbol{q}+\boldsymbol{Q}}(\boldsymbol{G})$, BSE eigenvector components $X_{v,c,\boldsymbol{q}}^{(m,\boldsymbol{Q})}$ and $GW$ energies $\varepsilon_{v,\boldsymbol{q}}$ to obtain the photoemission intensity $I_{m,\boldsymbol{Q}}(\boldsymbol{k})$, without the need for a BZ summation over $\boldsymbol{q}$. Given a sufficiently fine BZ sampling, the photoemission intensity for any $\boldsymbol{k}$ not on this grid (but within range of the energy cutoff used for wave functions in the DFT calculation) can then be obtained using efficient interpolation techniques.
Appendix C Computational Details
@@
For the example of monolayer hBN, we calculate Kohn-Sham DFT wave functions and energies in Quantum Espresso [43, 44, 45] using the Perdew-Burke-Ernzerhof exchange-correlation functional, ONCV pseudopotentials and a cutoff energy of 50 Ry and 200 Ry for the wave functions and the representation of the density, respectively. The lattice constant was set to 2.507 Å(bulk experimental value: 2.506 Å [55]) and we use a vacuum layer of 20 Å, as well as a truncation of the Coulomb interaction in $z$ direction [56], in order to avoid spurious interaction between repeated slabs from the periodic boundary condition. Based on these ground-state properties, we perform a $G_{0}W_{0}$ calculation using BerkeleyGW [1, 2, 47]. In particular, we use the NNS subsampling technique, which, for systems with reduced dimensionality, accelerates convergence with respect to the Brillouin-zone sampling [57]. For the calculation of the static inverse dielectric function we use a 12x12x1 sampling of the Brillouin-zone, a cutoff energy of 30 Ry for reciprocal lattice vectors and sum over 200 bands. Based on this, we calculate the $G_{0}W_{0}$ energies for 4 valence and 44 conduction bands, using the same k-space mesh and reciprocal lattice cutoff as for the inverse dielectric function.
@@
Finally, we construct the BSE Hamiltonian using 4 valence and 9 conduction bands on a 12x12x1 k-space mesh, interpolate it to a 96x96x1 k-space mesh (4 valence, 4 conduction bands) and diagonalize it.